# PUSSM: Point Cloud Upsampling as Implicit Statistical Shape Model

### v3

@@ Abstract
This paper proposes a framework for high-fidelity reconstruction of pelvic structures by integrating medical image segmentation and point cloud upsampling. By point cloud upsampling to learn shape priors from MedShapePelvic without requiring landmarks or PCA, our method functions as an implicit statistical shape model. Evaluations on Pelvic1k show significant improvements in surface quality and anatomical accuracy. This approach is generalizable and applicable to other skeletal regions.
Keywords:
Statistical Shape Modeling Point Cloud Upsampling Bone Shape Pelvic Segmentation.
1 Introduction
Statistical Shape Models (SSM) are based on traditional statistical methods designed to describe the average shape and density distribution, as well as the main variation patterns within these distributions. It is undeniable that U-Net[25] and its various deep learning segmentation models based on the U-Net architecture[6, 26] have consistently achieved state-of-the-art (SOTA) results in organ and bone segmentation tasks[36]. Additionally, various Segment Anything Model (SAM) approaches for medical imaging[15, 28, 21] are becoming increasingly prevalent. Despite this, SSM remains widely used across various domains, particularly in tasks such as image segmentation and measurement[2], and medical diagnosis[14]. Patient-specific models based on SSM are used in numerous medical applications, such as generating patient-specific bone models from sparse 3D data or 2D image sets using 3D SSM[27], or for in-vivo bone motion tracking[3].
Figure 1: The proposed Point Cloud Upsampling as Statistical Shape Model involves training a point cloud upsampling network to learn bone shape distributions from volumetric medical images. The image is then segmented into a point cloud, upsampled, and reconstructed into a complete model.
Creating ground truth for SSM differs from the masks used in semantic segmentation. Training an SSM requires manually identifying shape landmarks to infer statistical data. To build these shapes, a template grid is gradually deformed to segment training images and match the target boundaries, such as those of organs or bones.
To evaluate new patient scans using SSM, the new scan must be segmented and aligned, and the SSM must be optimized for the entire shape population. In this regard, Adams and Elhabian[1] introduced an unsupervised method for predicting the Point Distribution Model (PDM), enabling point cloud networks to learn accurate anatomical SSMs. Moreover, deep learning methods, inspired by renowned shape benchmarks in Computer Vision (CV) have further developed MedShapeNet[16], which is based on real patient imaging data and instruments. However, classical SSM construction typically involves manual annotation of anatomical landmarks and tedious correspondence alignment, which may limit scalability and automation in large datasets.
In contrast, our approach explores a data-driven alternative to SSM, where prior anatomical knowledge is learned implicitly through point cloud upsampling networks trained on large-scale 3D medical datasets, which is on the Pelvic model of MedShapeNet dataset (MedShapePelvic)[16] with the goal of learning realistic bone shape features through point cloud upsampling. Instead of modeling shape variation explicitly via PCA, our method leverages learned feature distributions from dense anatomical shapes to enhance the reconstruction of sparse pelvic scans. This shift from explicit modeling to learned priors enables more flexible and automatic shape analysis, especially when integrated with modern segmentation tools such as SAM-Med3D. This process transforms original, resolution-limited medical images and sparse voxels into denser, high-resolution 3D bone models, thereby improving shape reconstruction accuracy. The integration of medical image segmentation and point cloud upsampling within this framework enables SSM to describe the average shape and facilitates segmentation and model reconstruction.
In summary, the key contributions of our work are as follows:
1) We demonstrate the necessity of learning prior knowledge from MedShapePelvic by sampling the pelvic model as point clouds and constructing a baseline using four SOTA point cloud upsampling models, comparing the results with those based on PU1k.
2) We apply the SAM-Med3D model to the Pelvic1k subdataset to obtain predicted segmented voxels, convert the voxels into point clouds, and input them into the upsampling model. This process results in smoother, more complete 3D shape reconstructions.
3) The proposed framework, which integrates medical image segmentation and point cloud upsampling, has the potential to be transferred to various other skeletal models and organs.
2 Related Works
2.1 Point Cloud Processing
2.1.1 In 3D Object
@@
Since the introduction of PointNet[22, 23], this direct processing avoids the loss of context and complex steps involved in the data conversion process. Point cloud completion, in particular, is used to extract global features of 3D model, which can help supplement geometric details[18, 17]. With the aid of deep learning and large-scale 3D datasets, learning-based methods have achieved excellent performance in shape completion tasks. For example, in point cloud processing tasks, PCN[33] first employs a similar encoder to extract features and outputs dense and completed point clouds from sparse and incomplete input. Later, ASFM-Net[29] used parallel global and local feature matching to reasonably infer missing geometric details of objects.
@@
Besides, PU-Net[32] is a pioneering work that introduces CNNs for point cloud upsampling. Subsequently, PU-GCN[24] employed a graph-based network construction, achieving good upsampling performance. Grad-PU[13] similarly provides flexible upsampling at different scales. PUCRN[9] optimized point cloud generation and performance through a multi-stage structure and Transformers[35], especially in sparse or missing regions. However, these methods are often trained on synthetic datasets (e.g., PU1k dataset[24] proposed by PU-GCN) and may struggle to generalize to real anatomical shapes with irregular geometry and sparse sampling. Moreover, they typically focus on object-level CAD models and are not directly validated on medical segmentation pipelines.
2.1.2 In Medical Image
For medical image applications, Chen et al.[7] proposed a learning grid to point cloud deformable shape registration framework from multiple 2D image slices, which can achieve accurate 3D grid reconstruction in real-time in the absence of data and sparse annotations. Point cloud completion was first explored by Adams and Elhabian[1], who used [33] in an unsupervised manner to predict the Point Distribution Model (PDM) from 3D point clouds, exploring the potential of learning SSM from point clouds. And Gu et al. [10] proposed a new approach for 3D reconstruction of individual X-ray images, using point cloud completion to reconstruct real bones from unseen CT images. Later, Yassin et al.[30] proposed a multimodal foundation model for medical point cloud completion based on MedShapeNet[16]. This model is designed to handle incomplete 3D point cloud data and reconstruct the complete shape of various medical structures. By combining 3D point cloud data with textual data, the model improves the accuracy of shape reconstruction. Zhang et al.[34] proposed a hierarchical feature learning framework based on the State Space Model for medical point cloud understanding, which will aid in anatomical classification, completion, and segmentation.Although point cloud completion shares similarities with point cloud upsampling tasks by extracting global features from partial input to complement geometric details, no current work has applied these techniques to 3D point cloud data in the context of SSM to reconstruct various medical structures.
In contrast to these works, our approach explicitly integrates a medical image segmentation model (SAM-Med3D) with point cloud upsampling to form a complete pipeline from imaging to shape modeling. Furthermore, by training on realistic datasets, MedShapePelvic, and test on Pelvic1k subdataset, our method captures clinically relevant anatomical features and improves the quality of reconstructed surfaces.
2.2 Segment Anything
@@
3 Methodology
3.1 Overview of the Proposed Framework
@@
We propose a novel framework called Point Cloud Upsampling as Implicit Statistical Shape Model (PUSSM), which integrates medical image segmentation and point cloud upsampling for high-fidelity anatomical reconstruction. As shown in Figure 1(a), the core idea is to learn an implicit shape prior from real anatomical data through a deep upsampling network, allowing sparse point clouds to be restored to dense representations in an anatomically plausible way.
@@
The pipeline consists of two sequential modules. First, we apply the SAM-Med3D model to volumetric CT data to obtain voxel-wise segmentation masks. These masks are converted into sparse point clouds using spatial coordinates. Then, as illustrated in Figure 1(b), the sparse point cloud is passed into a pretrained point cloud upsampling network, which has been trained on MedShapePelvic samples. Finally, surface reconstruction is applied to the dense point cloud using Marching Cubes.
@@
Figure 2: Distributional interpretation of implicit statistical shape modeling. The segmentation-derived point cloud (left) deviates from the real anatomical distribution. Our network $f_{\theta}$ learns to map the sparse inputs toward the ground-truth distribution through point cloud upsampling.
3.2 Implicit Statistical Shape Modeling
Unlike classical Statistical Shape Models (SSMs), which explicitly encode anatomical variation via PCA over aligned meshes, our framework learns an implicit statistical shape model through data-driven training. Specifically, let $X\subset\mathbb{R}^{3}$ denote a sparse point cloud from SAM-Med3D, and let $Y\subset\mathbb{R}^{3}$ be the corresponding ground truth dense point cloud from MedShapePelvic.
We train a deep function $f_{\theta}:\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}^{M\times 3}$ to minimize the geometric discrepancy between $f_{\theta}(X)$ and $Y$, using a distance metric $D(\cdot,\cdot)$ such as the Chamfer Distance:
$\min_{\theta}\mathbb{E}_{(X,Y)\sim\mathcal{D}}\left[D(f_{\theta}(X),Y)\right]$
This process embeds the shape prior directly into the network weights $\theta$, enabling the model to generalize to new, noisy, or incomplete segmentation results. Thus, $f_{\theta}$ serves as an implicit statistical shape model learned directly from anatomical data—requiring no landmark alignment, shape vectors, or PCA decomposition.
@@ 3.2 Implicit Statistical Shape Modeling
Figure 2 further illustrates this concept from a distributional perspective, showing how our model maps the segmentation-derived distribution toward the ground-truth anatomical distribution via upsampling.
Figure 3: Conceptual comparison between classical SSM and our implicit PUSSM framework across key dimensions.
3.3 Comparison with Classical SSMs
To highlight the differences between our method and classical SSMs, Figure 3 provides a conceptual comparison across three key dimensions: shape representation, prior data source, and learning strategy. Traditional SSMs rely on manually constructed shape vectors and PCA decomposition of landmark-aligned meshes. In contrast, our method encodes shape priors directly through end-to-end learning from dense mesh data, enabling greater flexibility and broader applicability.
3.4 Evaluation Metrics
3.4.1 For Point Cloud Upsampling
@@
[…]
3.4.2 For Surface Reconstruction
@@
Area-Length Ratio (ALR) measures the shape quality of individual triangles within a mesh. This ratio attains its maximum value of 1 for equilateral triangles, indicating optimal shape quality. It is defined as:
Manifoldness Rate (MR) quantifies the proportion of edges in a mesh that satisfy manifold conditions. An edge is considered manifold if it is shared by exactly two faces. A rate approaching 1 indicates a well-structured mesh suitable for operations like Boolean computations and 3D printing. Conversely, a lower rate highlights the presence of non-manifold edges, which may necessitate mesh repair or refinement. It is defined as:
Manifoldness Rate=Number of manifold edgesTotal number of $\text{Manifoldness Rate}=\frac{\text{Number of manifold edges}}{\text{Total number of edges}}$
Connected Component Discrepancy (CC Diff.) assesses the topological connectedness of a mesh by comparing the observed number of connected components to the expected number. A discrepancy of 0 signifies that the mesh has the anticipated number of connected components, implying structural integrity. It is defined as:
[…]
@@
4 Experiments
4.1 Settings
4.1.1 Datasets
In our study, experimental validation on the pelvic region have been conducted in the hips, femurs, and sacrum from MedShapeNet[16] as MedShapePlevic for point cloud upsampling to learn shape prior knowledge. We then visualized Quantitative analyzed result based on Pelvic1k[19], a comprehensive pelvic CT dataset capable of replicating actual appearance changes. Due to the lack of maintenance in some of the open source data used by Pelvic1k, we can only obtain the following data (e.g. subdatasets 3, 4, 6).
For training and testing the upsampling network, we used the MedShapePelvic dataset based on MedShapeNet[16], which includes 50,496 training samples from 3,156 Mesh models generated via patch sampling[31]. The dataset consists of 3,506 3D models, split into 3,156 training samples and 350 testing samples. Patch sampling was performed by generating training data through Poisson disk sampling on patches of the 3D meshes.
4.1.2 Comparison Methods
For a fair and objective comparison, we obtained the open-source codes of these methods and trained and tested them on our computing equipment. And we test the result based on training in MedShapePelvic[16] and PU1k dataset[24], which is a synthetic dataset from 3D models. This comparison with identical parameter settings would be presented in ablation: Section 4.3.
4.1.3 Implementation Details
4.2 Upsampling Results
In accuracy analysis, as mentioned above, the generalization performance of two SOTA network have been evaluated on the MedShapePelvic dataset in 4×4$\times$4 × upsampling. The quantitative results for the dataset in Table 1 show that our method performs consistently across different input scales (512, 1024, 2,048 and 4,096 points) and outperforms the original network in CD, HD, and P2F metrics.
Besides, we also conducted 16×16$\times$16 × upsampling tests, in Table 2. Grad-PU can perform upsampling at any scale, so we set its ratio directly to 16×16$\times$16 × on the test set. However, the testing model of PUCRN can only use the same upsampling factor as the training conditions, so we performed two 4×4$\times$4 × samplings on PUCRN to achieve a 16×16$\times$16 × sampling effect. The results show that PUCRN still has good performance.
Table 1: Performance comparison of two methods with MedShapePelvic across input scales.
@@ 4.2 Upsampling Results
4.3 Ablation Studies
To verify the effectiveness and necessity of training in MedShapePelvic to learn the prior knowledge of shape, the 4×4$\times$4 × sampling result based on training in MedShapePelvic and PU1k dataset[24] have been tested. Among them, MedShapePelvic has 50496 training samples and PU1k has 69000 training samples. This avoids the difference in learning effectiveness caused by the large difference in the number of training samples. The result shows in Table 3
Table 2: Quantitative comparison in 16×16$\times$16 × upsampling after training in 4×4$\times$4 ×. The input points are 512 and 1024.

### v4

# Med-PU: Point Cloud Upsampling for High-Fidelity 3D Medical Shape Reconstruction

@@ Abstract
High-fidelity 3D anatomical reconstruction is a prerequisite for downstream clinical tasks such as preoperative planning, radiotherapy target delineation, and orthopedic implant design. We present Med-PU, a knowledge-driven framework that integrates volumetric medical image segmentation with point cloud upsampling for accurate pelvic shape reconstruction. Unlike landmark- or PCA-based statistical shape models, Med-PU learns an implicit anatomical prior directly from large-scale 3D shape data, enabling dense completion and refinement from sparse segmentation-derived point sets. The pipeline couples SAM-Med3D-based voxel segmentation, point extraction, deep upsampling, and surface reconstruction, yielding smooth and topologically consistent meshes. We evaluate Med-PU on pelvic CT datasets (MedShapePelvic for training and Pelvic1k for validation), benchmarking against state-of-the-art upsampling methods using comprehensive geometry and surface metrics. Med-PU consistently improves surface quality and anatomical fidelity while reducing artifacts, demonstrating robustness across input densities. Although validated on the pelvis, the approach is anatomy-agnostic and applicable to other skeletal regions and organs. These results suggest Med-PU as a practical, generalizable tool to bridge segmentation outputs and clinically usable 3D models.
I Introduction
High-fidelity 3D anatomical reconstruction is a prerequisite for several downstream clinical tasks, including preoperative planning, radiotherapy target delineation, and orthopedic implant design. In routine pipelines, clinicians often begin with voxel-wise segmentation and then derive surface meshes; however, directly converting voxel outputs into smooth, watertight, and anatomically faithful meshes remains challenging due to partial observations, image noise, and inter-subject variability.
Traditional Statistical Shape Models (SSMs) address variability by aligning training shapes and learning a low-dimensional subspace for population-level morphometrics. They have been widely adopted for segmentation, measurement, and diagnosis [1, 2], and can even generate patient-specific bone models from sparse 3D/2D data [3, 4]. Nevertheless, classical SSM construction depends on manual landmark annotation, explicit correspondence, and PCA, which limits scalability and automation on large medical datasets and can constrain the expressiveness of subtle anatomical details.
In parallel, modern segmentation networks—U-Net variants [5, 6, 7], nnFormer [8], and SAM-based extensions for medical imaging [9, 10, 11]—deliver robust voxel-level delineations. Yet a persistent gap remains between these voxel labels and clinically usable 3D models: meshes reconstructed from discrete masks are often jagged, topologically inconsistent, or over-smoothed, especially under limited resolution or incomplete coverage. This motivates learning shape priors directly from data so that sparse or noisy anatomical evidence can be lifted to dense, high-quality surfaces.
Recent studies are beginning to bridge point cloud learning and medical applications. Adams and Elhabian [12] explored unsupervised point-distribution modeling from raw point sets, while MedShapeNet [13] established a large-scale benchmark for medical 3D shape learning. These advances indicate that implicit, data-driven priors can complement or replace explicit SSMs in clinical 3D modeling, particularly when combined with strong segmentation front-ends.
Figure 1: Overview of the proposed pipeline. Medical images are first segmented into coarse anatomical masks (e.g., SAM-Med3D), then converted to point clouds, followed by upsampling to reconstruct a complete and high-resolution 3D bone model. Unlike classical Statistical Shape Models that require landmarks and PCA, our framework learns shape variability implicitly from dense point cloud data.
This work
We propose a data-driven framework for producing clinically amenable meshes from segmentation outputs. Instead of explicitly constructing and sampling a statistical shape space, we leverage point cloud upsampling networks trained on large-scale 3D medical shape data to implicitly encode anatomical priors. Using the pelvic subset of MedShapeNet (MedShapePelvic) as shape supervision and SAM-Med3D as a volumetric segmentation front-end, our framework learns to complete and refine sparse, segmentation-derived point sets into dense, anatomically faithful surfaces. This shift from explicit parametric modeling to learned priors enables flexible, automated reconstruction under varying input densities, helping to bridge the gap between voxel labels and high-quality meshes required in downstream clinical workflows.
Contributions
  • A segmentation-to-mesh pipeline with implicit priors. We integrate SAM-Med3D-based voxel segmentation with deep point cloud upsampling to learn an implicit anatomical prior that densifies and regularizes segmentation-derived point sets, producing smooth, watertight meshes without manual landmarks or PCA.
  • Comprehensive evaluation on pelvic CT. Trained on MedShapePelvic and validated on Pelvic1k, the proposed method is benchmarked against state-of-the-art upsampling approaches using geometry- and surface-level metrics (Chamfer/Hausdorff distances, point-to-surface error, F-score, normal consistency, and edge-aware measures), demonstrating consistent gains in surface quality and anatomical fidelity.
  • Generalization potential beyond the pelvis. Although evaluated on pelvic bones, the framework is anatomy-agnostic and applicable to other skeletal regions and organs, providing a practical bridge from voxel segmentation to clinically usable 3D models in broader biomedical contexts.
II Related Works
II-A Point Cloud Processing
II-A1 In 3D Object
@@
Since the introduction of PointNet[14, 15], direct processing of raw point sets avoids the loss of context and complex steps involved in voxelization or multi-view conversion. Point cloud completion, in particular, is used to extract global features of 3D models, which can help supplement geometric details[16, 17]. With the aid of deep learning and large-scale 3D datasets, learning-based methods have achieved excellent performance in shape completion tasks. For example, PCN[18] first employs an encoder to extract features and outputs dense and completed point clouds from sparse and incomplete input. Later, ASFM-Net[19] used parallel global and local feature matching to reasonably infer missing geometric details of objects.
@@
Besides, PU-Net[20] is a pioneering work that introduces CNNs for point cloud upsampling. Subsequently, PU-GCN[21] employed a graph-based network, achieving strong upsampling performance. Grad-PU[22] provides flexible upsampling at different scales, while PUCRN[23] optimized generation through multi-stage cascades. Transformer-based frameworks[24] also improved robustness in sparse or missing regions. However, these methods are often trained on synthetic datasets (e.g., PU1k[21]) and may struggle to generalize to real anatomical shapes with irregular geometry and sparse sampling. Moreover, they typically focus on object-level CAD models and are not directly validated on medical segmentation pipelines.
II-A2 In Medical Image
Recently, point cloud learning has been increasingly applied to medical image analysis. Chen et al.[25] proposed a learning-based deformable registration framework (MR-Net) to reconstruct 3D cardiac meshes from sparse 2D contours, showing real-time capability even with incomplete annotations. Beetz et al. developed Point2Mesh-Net, which combines point-based encoders with mesh-based decoders for cardiac surface reconstruction from cine MRI slices[26] , and later extended this line into multi-objective autoencoders for myocardial infarction prediction, enabling explainable outcome modeling from 3D cardiac shapes[27]. Hu et al.[28] introduced a GAN with tree-structured graph convolutions to generate 3D brain point clouds from a single 2D MRI slice. Jiang et al.[29] proposed CPAConv-POCO, which leverages continuous position-adaptive convolution for lung nodule 3D reconstruction. Chen et al.[30] applied completion networks to cartilage repair, using transformer-enhanced point cloud completion to recover femoral cartilage surfaces. For anatomical organ-level segmentation, Zhang et al.[31, 32] designed point-voxel fusion networks to segment Couinaud liver segments, showing superior performance on LiTS and 3Dircadb. More recently, Yassin et al.[33] proposed a multimodal foundation model for medical point cloud completion based on the large-scale MedShapeNet dataset[13], achieving generalizable completion across 240+ anatomical classes. Zhang et al.[34] introduced a hierarchical feature learning framework based on state-space models to further improve anatomical understanding tasks such as completion and segmentation. Although point cloud completion shares similarities with upsampling by extracting global features from partial input to complement geometric details, most existing methods are validated on synthetic or restricted datasets. In contrast, our approach explicitly integrates a medical image segmentation model (SAM-Med3D) with point cloud upsampling to form a complete pipeline from imaging to shape modeling. Furthermore, by training on realistic datasets such as MedShapePelvic and evaluating on the Pelvic1k subset, our method captures clinically relevant anatomical features and improves the quality of reconstructed surfaces.
II-B Segment Anything
@@
III Methodology
III-A Overview of the Proposed Framework
@@
We propose a novel framework that integrates medical image segmentation and point cloud upsampling for high-fidelity anatomical reconstruction. As shown in Figure 1(a), the core idea is to learn an implicit anatomical prior from real data through a deep upsampling network, allowing sparse point clouds to be restored to dense and anatomically plausible representations.
@@
The pipeline consists of two sequential modules. First, we apply the SAM-Med3D model to volumetric CT data to obtain voxel-wise segmentation masks. These masks are converted into sparse point clouds using spatial coordinates. Then, as illustrated in Figure 1(b), the sparse point cloud is passed into a pretrained point cloud upsampling network, which has been trained on MedShapePelvic samples. Finally, surface reconstruction is applied to the dense point cloud using Marching Cubes, yielding smooth triangular meshes suitable for quantitative evaluation and visualization.
@@
Figure 2: Distributional interpretation of implicit anatomical prioring. The segmentation-derived point cloud (left) deviates from the real anatomical distribution. Our network $f_{\theta}$ learns to map the sparse inputs toward the ground-truth distribution through point cloud upsampling. This reflects an implicit learning of intra-population variation, capturing not only shape completion but also a shift toward the anatomical manifold.
III-B Learning Implicit Anatomical Priors
Distributional notation
Let $X=\{x_{i}\}_{i=1}^{N}\subset\mathbb{R}^{3}$ be the sparse points sampled from SAM-Med3D masks and $Y=\{y_{j}\}_{j=1}^{M}$ the dense target points from MedShapePelvic. We regard point clouds as empirical measures $\mu_{X}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$ and $\nu_{Y}=\frac{1}{M}\sum_{j=1}^{M}\delta_{y_{j}}$. The upsampler $f_{\theta}:\mathbb{R}^{N\times 3}\!\to\!\mathbb{R}^{M\times 3}$ induces a pushforward measure $f_{\theta:}\mu_{X}=\frac{1}{M}\sum_{j=1}^{M}\delta_{\hat{y}_{j}}$ with $\hat{Y}=f_{\theta}(X)=\{\hat{y}_{j}\}_{j=1}^{M}$. Learning then amounts to aligning $f_{\theta:}\mu_{X}$ with $\nu_{Y}$ on the anatomical shape manifold approximated by the training set.
Empirical risk
We minimize a symmetric Chamfer discrepancy between the predicted and target point sets: Eq. (1)
$\min_{\theta}\;\mathbb{E}_{(X,Y)\sim\mathcal{D}}\Big[\underbrace{\int\!\!\min_{y\in Y}\|z-y\|_{2}^{2}\,\mathrm{d}\big(f_{\theta:}\mu_{X}\big)(z)}_{\text{pred}\to\text{target}}\;+\;\underbrace{\int\!\!\min_{\hat{y}\in\hat{Y}}\|z-\hat{y}\|_{2}^{2}\,\mathrm{d}\nu_{Y}(z)}_{\text{target}\to\text{pred}}\Big],$ (1)
which is exactly the usual (squared) Chamfer distance written in a distributional form. In practice, Eq. (1) reduces to
$\mathrm{CD}(\hat{Y},Y)=\frac{1}{M}\sum_{\hat{y}\in\hat{Y}}\min_{y\in Y}\|\hat{y}-y\|_{2}^{2}+\frac{1}{M}\sum_{y\in Y}\min_{\hat{y}\in\hat{Y}}\|y-\hat{y}\|_{2}^{2}.$
Optimizing Eq. (1) embeds shape regularity and population variability into the network parameters $\theta$, enabling robust densification of sparse, segmentation-derived inputs without landmarks, explicit correspondences, or PCA.
@@ III-B Learning Implicit Anatomical Priors
Figure 2 further illustrates this concept from a distributional perspective, showing how our model maps the segmentation-derived distribution toward the ground-truth anatomical distribution via upsampling. We observe a consistent reduction in Chamfer Distance variance across test samples after upsampling, indicating convergence to a learned anatomical distribution.
Ergo, the network aims to reduce geometric variance and improving reconstruction fidelity across test cases.
III-C Evaluation Metrics
III-C1 For Point Cloud Upsampling
@@
$\text{P2F}(P,S)=\frac{1}{|P|}\sum_{p\in P}\min_{s\in S}\|p-s\|$
III-C2 For Surface Reconstruction
@@
Area-Length Ratio (ALR) measures the shape quality of individual triangles within a mesh. This ratio attains its maximum value of 1 for equilateral triangles, indicating optimal shape quality.
Manifoldness Rate (MR) quantifies the proportion of edges in a mesh that satisfy manifold conditions. An edge is considered manifold if it is shared by exactly two faces. A rate approaching 1 indicates a well-structured mesh suitable for operations like Boolean computations and 3D printing. Conversely, a lower rate highlights the presence of non-manifold edges, which may necessitate mesh repair or refinement.
Connected Component Discrepancy (CC Diff.) assesses the topological connectedness of a mesh by comparing the observed number of connected components to the expected number. A discrepancy of 0 signifies that the mesh has the anticipated number of connected components, implying structural integrity.
@@
TABLE I: Performance comparison of two methods with MedShapePelvic across input scales.
@@
IV Experiments
IV-A Settings
IV-A1 Datasets
In our study, experimental validation on the pelvic region have been conducted in the hips, femurs, and sacrum from MedShapeNet[13] as MedShapePelvic for point cloud upsampling to learn shape prior knowledge. We then visualized Quantitative analyzed result based on Pelvic1k[39], a comprehensive pelvic CT dataset capable of replicating actual appearance changes. Due to the lack of maintenance in some of the open source data used by Pelvic1k, we can only obtain the following data (e.g. subdatasets 3, 4, 6).
For training and testing the upsampling network, we used the MedShapePelvic dataset based on MedShapeNet[13], which includes 50,496 training samples from 3,156 Mesh models generated via patch sampling[40, 21, 41, 42]. The dataset consists of 3,506 3D models, split into 3,156 training samples and 350 testing samples. Patch sampling was performed by generating training data through Poisson disk sampling on patches of the 3D meshes.
IV-A2 Comparison Methods
For a fair and objective comparison, we obtained the open-source codes of these methods and trained and tested them on our computing equipment. And we test the result based on training in MedShapePelvic[13] and PU1k dataset[21], which is a synthetic dataset from 3D models. This comparison with identical parameter settings would be presented in ablation: Section IV-C.
IV-A3 Implementation Details
TABLE II: Quantitative comparison in $16\times$ upsampling after training in $4\times$. The input points are 512 and 1024.